OpenStudy (anonymous):

find three geometric means between 64 and 4

OpenStudy (sidsiddhartha):

4_ _ _ 64 64 / 4 = 16. Find its 4th root because there are 4 spaces. so \[\sqrt[4]{16}=2\] so second no is 2*4=8 third no 8*2=16 fourth no is 16*2=32 so three GMs are 8,16,32

OpenStudy (anonymous):

Thank you!

OpenStudy (sidsiddhartha):

wlcm :)